# Artificial Data Point Generation in Clustered Latent Space for Small Medical Datasets

### v1

@@ V-C Classification
Rage Baseline 54.00 ($\pm$39.80) 69.23 52.00 ($\pm$38.68) 53.84 64.00 ($\pm$38.78) 46.15 66.00 ($\pm$33.82) 53.84
GN 58.00 ($\pm$39.19) 69.23 56.00 ($\pm$38.26) 53.84 68.00 ($\pm$37.09) 46.15 72.00 ($\pm$31.87) 46.15
kNNMTD 54.00 ($\pm$39.80) 69.23 56.00 ($\pm$38.26) 53.84 64.00 ($\pm$38.78) 53.84 54.00 ($\pm$42.24) 61.53
AGCL 60.00 ($\pm$40.00) 76.92 58.00 ($\pm$41.67) 69.23 68.00 ($\pm$39.70) 53.84 74.00 ($\pm$32.00) 53.84
Happy Baseline 70.83 ($\pm$40.61) 61.11 70.83 ($\pm$43.10) 72.22 79.17 ($\pm$35.11) 66.66 75.00 ($\pm$38.19) 66.67
GN 70.83 ($\pm$40.61) 61.11 79.17 ($\pm$35.11) 72.22 79.17 ($\pm$35.11) 66.66 81.25 ($\pm$31.66) 77.78
kNNMTD 70.83 ($\pm$40.61) 61.11 79.17 ($\pm$37.96) 72.22 81.25 ($\pm$34.80) 66.68 68.75 ($\pm$42.85) 61.11
AGCL 70.83 ($\pm$40.61) 61.11 79.17 ($\pm$35.11) 83.33 81.25 ($\pm$34.80) 72.22 81.25 ($\pm$31.66) 83.33
Disgust Baseline 68.00 ($\pm$37.09) 46.15 76.00 ($\pm$32.00) 46.15 78.00 ($\pm$28.57) 61.54 70.00 ($\pm$40.00) 69.23
GN 76.00 ($\pm$34.99) 53.85 80.00 ($\pm$31.62) 38.46 82.00 ($\pm$24.00) 69.23 82.00 ($\pm$27.86) 30.77
kNNMTD 68.00 ($\pm$37.09) 46.15 76.00 ($\pm$34.99) 46.15 84.00 ($\pm$23.32) 61.54 72.00 ($\pm$37.63) 61.54
AGCL 72.00 ($\pm$37.63) 69.23 82.00 ($\pm$31.24) 53.84 84.00 ($\pm$30.72) 69.23 84.00 ($\pm$27.28) 76.92
Fear Baseline 60.00 ($\pm$40.00) 69.23 68.00 ($\pm$37.09) 84.62 62.00 ($\pm$38.16) 61.53 62.00 ($\pm$38.16) 53.84
GN 64.00 ($\pm$36.11) 69.23 66.00 ($\pm$39.29) 76.92 66.00 ($\pm$36.66) 61.53 70.00 ($\pm$37.42) 76.92
kNNMTD 60.00 ($\pm$40.00) 69.23 68.00 ($\pm$39.70) 84.62 72.00 ($\pm$31.87) 53.84 68.00 ($\pm$39.70) 61.53
AGCL 62.00 ($\pm$38.16) 76.92 70.00 ($\pm$40.00) 84.62 74.00 ($\pm$37.74) 69.23 72.00 ($\pm$37.63) 92.30
@@
GN 72.00 ($\pm$37.63) 50.00 70.00 ($\pm$37.42) 43.75 80.00 ($\pm$31.62) 50.00 76.00 ($\pm$34.99) 43.75
kNNMTD 68.00 ($\pm$39.70) 56.25 68.00 ($\pm$39.70) 50.00 68.00 ($\pm$34.29) 50.00 72.00 ($\pm$37.63) 56.25
AGCL 70.00 ($\pm$37.42) 62.50 78.00 ($\pm$34.87) 50.00 82.00 ($\pm$31.24) 56.25 84.00 ($\pm$30.72) 75.00
Voting Baseline 68.18 61.11 75.00 66.66 77.27 61.11 65.90 72.22
GN 70.45 66.66 70.45 61.11 81.81 66.66 77.27 66.66
kNNMTD 68.18 61.11 77.27 66.66 79.54 66.66 77.27 66.66
AGCL 75.00 77.77 79.54 77.77 90.90 77.77 84.09 83.33
In the previous section, we analyzed the data statistically. In this section, we evaluate how the data behaves when subjected to machine learning algorithms for classifying Parkinson’s and control data. A two-leave-out cross-validation method was applied to the training data, resulting in 25 unique folds. Each fold served as an independent subset for training separate models using the same hyperparameters. Throughout the cross-validation process, hyperparameter tuning was carried out to identify the optimal hyperparameters based on performance across all folds. The test data consisted of 11 control samples and 7 Parkinson’s disease samples. Instead of selecting a single model, all 25 models generated during the cross-validation process were applied to the test data. Their predictions were aggregated using a weighted majority voting approach to produce the final classification.
By comparing the effectiveness of AGCL against baseline, perturbation mechanisms and kNNMTD [16], we aim to evaluate its performance in classifying facial expressions. The baseline approach serves as a reference point without any augmentation, providing a foundational measure of classification performance.
In contrast, the perturbation-based synthetic data generation mechanism is stated as: $x^{s}=x_{i}+\epsilon$ where $\epsilon$ is a noise sampled from a certain distribution. According to Gaussian approach, $\epsilon$ is sampled with $N(0,\sigma)$. This method usually utilizes n-way marginals for purposes of creating variability in the categorical features [31]. Therefore, we employed the Gaussian approach to add noise to each sample in the latent space. To elucidate the impact of different levels of noise, we consider a range of standard deviations ($\sigma$= 0.001, 0.005, 0.01, 0.025, 0.05, 0.075, 0.1, 0.2, 0.3, 0.4, 0.5, 0.6, 0.8, 1]) to identify the optimal performance of Gaussian Noise (GN) for comparison with the AGCL method.
Moreover, We estimated the number of generated data points in the cluster for AGCL through grid search, gradually testing the $\alpha$ value from 2 to 60 with a step size of 5.
Table IV shows the performance of algorithms across four machine learning classifiers in screening Parkinson across five emotions: Rage, Happy, Disgust, Fear, and Surprise. To aggregate results obtained from five facial expressions, we employ the majority voting integration method, by voting from all facial expressions for each individual. This ensemble method helps in reducing the possibility of bias or errors that come with separate facial expressions, making it a more accurate and diverse approach. The metrics reported include both cross-validation and test accuracy.
For the rage emotion, AGCL demonstrated superior performance compared to other methods across all classifiers. The best validation accuracy was achieved using the MLP classifier, where AGCL reached 74.00%, whereas the baseline validation accuracy was lower at 66.00%. In the test results, AGCL also led with the highest accuracy of 76.92% with kNN, outperforming the other algorithms test accuracy of 69.23%.
In the happy emotion, AGCL achieved the highest validation accuracy of 81.25% with MLP, significantly improving over the kNNMTD’s validation accuracy of 68.75%. The test accuracy of 83.33% with both SVM and MLP shows the effectiveness of AGCL in generating high-quality synthetic data that enhances classification performance, compared to the GN and kNNMTD. While AGCL performed well in test accuracy compared to the baseline and other augmentation methods, the overall improvement in cross-validation was less pronounced than in other emotions. The validation accuracy of 81.25% with RF and MLP showed moderate gains, which suggests that the happy expression might be easier for classifiers to distinguish even without extensive data augmentation.
For the disgust emotion, AGCL also led to noticeable performance improvements. The best validation accuracy of 84.00% was achieved using the Random Forest classifier. AGCL also reached a test accuracy of 76.92% with MLP, which is the best test accuracy across all four algorithms. The fear emotion further highlights the advantages of AGCL in small dataset classification. AGCL achieved the highest validation accuracy of 74.00% with the Random Forest classifier, while the GN struggled with only 66.00%. On the test set, AGCL achieved its best accuracy with MLP, reaching an outstanding 92.30%, far exceeding the other algorithms.
In the surprise emotion, AGCL continued to outperform the baseline and other methods. The highest validation accuracy was 84.00% with MLP for AGCL. In the test phase, AGCL again demonstrated strong performance, achieving 75.00% accuracy with MLP, surpassing the baseline’s test accuracy of 62.50%. When results across all emotions were aggregated using majority voting, AGCL achieved the highest overall test accuracy of 83.33% with the MLP classifier. This marks a significant improvement over the other algorithms. The validation accuracy of 90.90% with Random Forest and 84.09% with MLP further confirms AGCL’s effectiveness in augmenting small medical datasets and improving classification results.
@@ VI Conclusion
In this paper, we introduced the Artificial Data Point Generation in Clustered Latent Space (AGCL) method to address the challenges of limited training data in medical applications, particularly in the classification of Parkinson’s disease through facial expression analysis. By generating synthetic data points in a clustered latent space, AGCL enhances the performance of machine learning models in scenarios with small datasets. The framework employs feature extraction, K-means clustering, and a separation criterion to guide synthetic data generation, ensuring that only well-separated clusters contribute to data augmentation.
Experimental results across multiple classifiers and emotions demonstrate that AGCL consistently outperforms baseline methods, as well as other data augmentation techniques like Gaussian noise and kNNMTD. AGCL achieved the highest overall test accuracy of 83.33% through majority voting, confirming its effectiveness in improving classification in small datasets.
AGCL’s ability to generate synthetic data points tailored to the structure of the dataset offers significant advantages in enhancing the generalization capabilities of machine learning models. This method holds promise for broader applications in medical domains and other fields where data scarcity is a critical issue. Future work may explore the extension of AGCL to different datasets and further refinement of clustering techniques to increase its applicability across a wider range of tasks.

### v2

@@ V-D Evaluation of Feature Space Augmentation Strategies
Rage Baseline $54.00(\pm 39.80)$ $69.23$ $52.00(\pm 38.68)$ $53.84$ $64.00(\pm 38.78)$ $46.15$ $54.00(\pm 42.24)$ $76.92$
GN $58.00(\pm 39.19)$ $69.23$ $56.00(\pm 38.26)$ $53.84$ $64.00(\pm 37.09)$ $46.15$ $64.00(\pm 32.00)$ $76.92$
kNNMTD $54.00(\pm 39.80)$ $69.23$ $56.00(\pm 38.26)$ $53.84$ $64.00(\pm 38.78)$ $53.84$ $54.00(\pm 39.80)$ $69.23$
FICAug $60.00(\pm 40.00)$ $76.92$ $58.00(\pm 41.67)$ $69.23$ $68.00(\pm 39.70)$ $53.84$ $64.00(\pm 38.78)$ $84.61$
Happiness Baseline $70.83(\pm 40.61)$ $61.11$ $70.83(\pm 43.10)$ $72.22$ $79.17(\pm 35.11)$ $66.66$ $75.00(\pm 38.19)$ $77.77$
GN $70.83(\pm 40.61)$ $61.11$ $70.83(\pm 35.11)$ $72.22$ $79.17(\pm 35.11)$ $66.66$ $79.17(\pm 35.11)$ $66.66$
kNNMTD $70.83(\pm 40.61)$ $61.11$ $70.83(\pm 37.96)$ $72.22$ $79.17(\pm 34.80)$ $66.66$ $68.75(\pm 42.85)$ $72.22$
FICAug $70.83(\pm 40.61)$ $61.11$ $79.17(\pm 35.11)$ $83.33$ $81.25(\pm 34.80)$ $72.22$ $81.25(\pm 34.80)$ $77.77$
Disgust Baseline $68.00(\pm 37.09)$ $42.85$ $76.00(\pm 32.00)$ $42.85$ $78.00(\pm 28.57)$ $57.14$ $70.00(\pm 40.00)$ $57.14$
GN $76.00(\pm 34.99)$ $50.00$ $80.00(\pm 31.62)$ $35.71$ $82.00(\pm 24.00)$ $64.28$ $80.00(\pm 31.62)$ $35.71$
kNNMTD $68.00(\pm 37.09)$ $42.85$ $76.00(\pm 34.99)$ $42.85$ $84.00(\pm 23.32)$ $57.14$ $68.00(\pm 42.14)$ $50.0$
FICAug $72.00(\pm 37.63)$ $64.28$ $82.00(\pm 31.24)$ $50.00$ $84.00(\pm 30.72)$ $64.28$ $82.00(\pm 27.86)$ $64.28$
Fear Baseline $60.00(\pm 40.00)$ $69.23$ $68.00(\pm 37.09)$ $84.62$ $58.00(\pm 39.19)$ $69.23$ $62.00(\pm 40.69)$ $61.53$
GN $64.00(\pm 41.28)$ $69.23$ $68.00(\pm 39.70)$ $76.92$ $64.00(\pm 43.63)$ $69.23$ $70.00(\pm 37.42)$ $69.23$
kNNMTD $60.00(\pm 40.00)$ $69.23$ $68.00(\pm 39.70)$ $76.92$ $72.00(\pm 31.87)$ $53.84$ $72.00(\pm 37.63)$ $61.53$
FICAug $62.00(\pm 43.08)$ $69.23$ $72.00(\pm 37.63)$ $76.92$ $74.00(\pm 34.99)$ $69.23$ $76.00(\pm 34.99)$ $69.23$
@@
GN $68.00(\pm 37.63)$ $50.00$ $70.00(\pm 37.42)$ $43.75$ $80.00(\pm 31.62)$ $50.00$ $78.00(\pm 31.87)$ $43.75$
kNNMTD $68.00(\pm 39.70)$ $56.25$ $68.00(\pm 39.70)$ $50.00$ $68.00(\pm 34.29)$ $50.00$ $74.00(\pm 34.99)$ $62.50$
FICAug $70.00(\pm 37.42)$ $62.50$ $78.00(\pm 34.87)$ $50.00$ $82.00(\pm 31.24)$ $56.25$ $82.00(\pm 31.24)$ $75.00$
All Emotions Baseline $68.18$ $61.11$ $72.72$ $66.66$ $75.00$ $66.66$ $68.18$ $72.22$
GN $70.45$ $66.66$ $72.72$ $61.11$ $82.36$ $66.66$ $81.81$ $66.66$
kNNMTD $68.18$ $61.11$ $77.27$ $66.66$ $79.54$ $66.66$ $75.00$ $77.77$
FICAug $75.00$ $72.22$ $79.54$ $72.22$ $84.09$ $77.77$ $81.81$ $83.33$
V-E GANimation and Synthetic Face Generation
To initiate the generation process in the original image space, we first utilized the Generated Photos tool [36] to create a collection of neutral facial images. This tool employs GAN-based models to synthesize realistic human faces. To increase the realism and relevance of the dataset, we specifically selected images of elderly individuals with neutral expressions (i.e., without any emotional cues). These images provided a strong foundation for applying the synthetic AUs. Examples of these synthesized neutral faces are shown in Figure 3.
Figure 3: AI-generated neutral face images used for applying Action Units
The GANimation  [ganimation] model is a generative adversarial network specifically designed for facial expression synthesis. It allows for precise and continuous control over facial movements by applying AUs to a given face, without requiring paired training data. Unlike traditional conditional GANs, GANimation uses attention and deformation mechanisms to focus only on the facial regions that need to be modified, leading to more natural and expressive outputs. Its architecture is particularly well-suited for generating identity-preserving facial expressions across various emotional states.
To improve the performance and accuracy of the GANimation model, we fine-tuned it using the generated neutral images. The model was fine-tuned for 20 epochs on this dataset, allowing it to adapt to the specific characteristics of our data and learn relevant facial features.
In the next step, we fed the GANimation model with the synthesized neutral images along with facial AUs generated for each class using the clustering algorithm. These AUs were created in two ways: (1) a set of dominant AUs selected for each emotion based on our proposed method involving clustering, cluster evaluation, and re-clustering, and (2) additional AUs sampled from a normal distribution to increase diversity and realism. The normal distribution helped expand the variation in generated expressions and improve the representation of emotional diversity.
Finally, GANimation applied the given AUs to the neutral faces and produced diverse synthetic facial expressions for each class in the dataset.
The neutral faces used for both control and Parkinson’s groups were identical; the only difference lay in the AUs applied to them. Examples of the generated images for each emotion and both groups are illustrated in Figures 4 and 5.
Figure 4: Synthetic facial expressions generated for control individuals using GANimation
Figure 5: Synthetic facial expressions generated for Parkinson’s patients using GANimation
To evaluate the quality of the generated images, we used the OpenFace toolkit to extract AUs from the synthetic images. These extracted AU values were then compared to the input AUs fed to the GANimation model, and the Mean Absolute Error(MAE) was calculated. Table V presents the MAE values for each emotion and both subject groups. The low error values confirm the model’s ability to accurately reproduce the intended AUs.
TABLE V: Evaluation of GANimation using MAE
Emotion Rage Happiness Disgust Fear Surprise
Parkinson’s 0.17 0.09 0.08 0.19 0.09
Control 0.16 0.05 0.07 0.06 0.07
V-F Training the Convolutional Neural Network Model
To train the proposed model, we first used ResNet18 and trained it on the synthetic emotional face images. Since the neutral faces were identical for both groups, this initial training phase encouraged the model to focus on differences in emotional expression and Action Unit patterns, rather than on irrelevant visual differences between individual faces. After this pretraining phase, the model was fine-tuned on the original dataset.
Throughout the training process, various data augmentation techniques were employed to improve the model’s performance. During the first phase, when training on synthetic images, transformations such as horizontal flipping, rotation, and perspective distortion were applied. In the second phase, where the model was fine-tuned on the original dataset, Gaussian noise was additionally introduced. These augmentation strategies helped increase the diversity of training samples and enhanced the model’s generalization capability.
Table VI presents the classification results of FICAug compared to three baselines: a standard ResNet18, a DINO-pretrained ResNet50, and the Masked Siamese Networks(MSN) model using facebook/vit-msn-small. All models were trained under identical settings, with consistent data splits, augmentation techniques, training epochs, and optimizer configurations. This ensures a fair comparison between the methods.
Across all five emotions as well as the aggregated All Emotions category, FICAug consistently outperforms the baselines in both validation and test accuracy. The relatively high standard deviation observed in validation results stems from the use of a two-leave-out cross-validation strategy. Since each fold contains only two samples for validation (one from each class), any misclassification significantly affects the accuracy, leading to larger variance across folds.
TABLE VI: Comparison of classification accuracy between FICAug and other algorithms
Emotion Method Cross Validation Test
Rage Baseline $62.00(\pm 38.15)$ $53.84$
DINO $44.00(\pm 29.00)$ $30.76$
MSN $70.0(\pm 37.41)$ $61.53$
Proposed $74.00(\pm 37.73)$ $69.23$
Happiness Baseline $72.91(\pm 35.29)$ $72.22$
DINO $60.41(\pm 40.77)$ $55.55$
MSN $83.33(\pm 27.63)$ $55.55$
FICAug $81.25(\pm 34.80)$ $77.77$
Disgust Baseline $82.00(\pm 27.27)$ $71.42$
DINO $62.00(\pm 35.44)$ $50.00$
MSN $82.00(\pm 27.85)$ $64.28$
FICAug $86.00(\pm 26.53)$ $78.57$
Fear Baseline $70.00(\pm 34.64)$ $69.23$
DINO $52.00(\pm 29.93)$ $30.76$
MSN $76.00(\pm 32.00)$ $69.23$
FICAug $78.00(\pm 34.87)$ $76.92$
Surprise Baseline $68.00(\pm 37.09)$ $75.00$
DINO $40.00(\pm 31.62)$ $56.25$
MSN $80.00(\pm 34.64)$ $62.50$
FICAug $68.00(\pm 37.09)$ $75.00$
All Emotions Baseline $70.96$ $72.22$
DINO $50.00$ $66.66$
MSN $68.18$ $72.22$
FICAug $88.63$ $94.00$
In the Rage emotion, FICAug achieves a test accuracy of 69.23%, surpassing the baseline ResNet18 (53.84%), DINO (30.76%), and MSN (61.53%). For Happiness, although the MSN model showed the highest cross-validation accuracy (83.33%), its test performance dropped significantly to 55.55%. FICAug, in contrast, maintained strong generalization with a test accuracy of 77.77%.
In the case of Disgust, FICAug achieved the highest cross-validation accuracy (86.00%) and test accuracy (78.57%), outperforming all three baselines. Moreover, for the Fear category, our method again demonstrated superior performance, achieving the best test accuracy of 76.92%, compared to 69.23% for both the baseline and MSN, and only 30.76% for DINO.
In the Surprise category, FICAug matched the baseline with a test accuracy of 75.00%, while both DINO and MSN underperformed at 56.25% and 62.50%, respectively.
Aggregating results across all emotions, FICAug achieved a substantial improvement, reaching a test accuracy of 94.00%, compared to 72.22% for both the baseline and MSN, and 66.66% for DINO. The corresponding validation accuracy was also the highest at 88.63%, further confirming the generalization capabilities of our approach.
These results clearly demonstrate that FICAug, which combines feature-space clustering and synthetic image generation, is an effective strategy for enriching training data and improving classification performance—particularly in small data regimes. Unlike self-supervised approaches, our method generates new training samples by mapping existing, clustered latent features back to the image space, guided by an augmentation strategy designed to increase diversity and coverage of expression patterns. By training a CNN on these generated images, the model is not simply learning from duplicated or noised data—it is exposed to novel expressions synthesized from meaningful variations in the original features. This enables the extraction of new, discriminative representations that better reflect subtle emotional cues, as evidenced by the method’s superior performance.
V-G Comparison of Feature Quality: Feature Space vs. Image-Based Learning
In the final analysis, we compare classification performance between two strategies: direct classification using augmented latent features versus training a CNN on synthetic facial images reconstructed from those features. As shown in Figure 6, learning from image-based data using CNN leads to a substantial performance boost. The validation accuracy increased to 88.63%, compared to 81.81% in the feature space approach. Similarly, test accuracy improved markedly, reaching 94.00%, as opposed to 83.33% with feature space classification.
This improvement highlights the strength of our approach: by generating synthetic images from existing latent features and training a CNN on them, the model is able to extract new, richer representations—effectively learning new features from data derived from itself. This self-guided learning process enables the network to go beyond the original feature space and develop more expressive and discriminative patterns, ultimately improving accuracy.
Figure 6: Comparison of validation and test accuracy between latent space classification and image-based CNN training across all emotions.
@@ VI Conclusion
This study presented FICAug, a feature-to-image data augmentation framework aimed at improving machine learning performance under small-data conditions. The method first generates synthetic samples in the feature space using k-means clustering and Gaussian sampling. To avoid increasing data complexity, data are not generated in regions where class overlap exists. This targeted augmentation improves diversity while maintaining the structural integrity of the original distribution. Experimental evaluations on a facial expression dataset for Parkinson’s disease screening demonstrated that the generated data preserved key statistical properties of the original samples and significantly improved classification accuracy across multiple classifiers.
To enhance representation learning, the synthetic features were mapped back into the image space using a GANimation-based generative model. Notably, FICAug achieved a cross-validation accuracy of 88.63%, significantly surpassing the best baseline performance of 70.96%. This highlights the ability of FICAug to extract richer and more discriminative features from structured synthetic data.